OpenStudy (anonymous):

Does the equation represent a direct variation? If so, find the constant of variation. 5x = -3y A. no B. yes; k = -5/3 C. yes; k = 5/3 D. yes; k = 3/5

OpenStudy (anonymous):

direction variation is saying y varies directly as x. so they increase together. make the equation resemble y\[\frac{ y }{ x }=k\]

OpenStudy (anonymous):

\[\frac{ y }{ x }=k\] \[k =\frac{ -3 }{ 5 }\] so it's A?

OpenStudy (anonymous):

divide both sides by "x" from there multiply both sides by \[\ -\frac{ 1 }{ 3 }\] cool?

OpenStudy (anonymous):

no its not A